OpenStudy (wade123):

@satellite73 @Loser66

OpenStudy (wade123):

g(x) would grow faster than f(x) right?

OpenStudy (anonymous):

It will most of the time, but what about between 0 and 1?

OpenStudy (misty1212):

??

OpenStudy (misty1212):

a cube grows much faster than a square , although as @epiphi pointed out, not between 0 and 1, but after one it certainly does

OpenStudy (wade123):

ohh so it would be the other way around?

OpenStudy (wade123):

@Loser66

OpenStudy (loser66):

let f1= x^2 f2 = x^3 +1 then f'1= 2x f'2= 3x^2 Between the two, f1' and f2' which one is bigger?

OpenStudy (wade123):

f2'

OpenStudy (loser66):

sure, and f' is rate of change, dat sit

OpenStudy (wade123):

wait im confused hahaha @Loser66

OpenStudy (wade123):

so does g(x) grow fast or does f(x)